# Learning Part-aware 3D Representations by Fusing 2D Gaussians and Superquadrics

### v1

@@ IV-C2 ShapeNet Benchmark
Complete model $2.32$ $23.91$ $0.891$ $0.126$ $\underline{0.94}$ $\underline{36.07}$ $0.977$ $0.038$ $6.6$
w/o $L_{\text{over}}$ $\underline{2.11}$ $\underline{24.13}$ $\underline{0.896}$ $\underline{0.122}$ $1.34$ $36.04$ $\underline{0.977}$ $\underline{0.037}$ $7.6$
w/o $L_{\text{opa}}$ $5.72$ $20.12$ $0.853$ $0.174$ $1.06$ $35.60$ $0.975$ $0.040$ $3.8$
@@
w/o $L_{\text{par}}$ 1.91 24.47 0.898 0.119 0.93 36.18 0.977 0.037 $10.1$
IV-C3 Real-life Data
To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from BlenderMVS dataset [56] and a self-captured dataset. As shown in Fig. 5, our approach can robustly produce both realistic appearances and meaningful 3D decompositions across a variety of data types. Additionally, the distinct advantages of block-level and point-level reconstructions are clearly observable. For instance, in the sixth column, due to limited training views, point-level reconstructions based on 2DGS tend to generate artifacts, whereas block-level optimization produces accurate and smooth surfaces. Conversely, point-level reconstruction demonstrates superior capability in capturing intricate geometric details.
Figure 7: Impact of Gaussian scale regularization (SR). This figure demonstrates the effect of Gaussian scale regularization. As shown by the typically degraded mesh reconstruction from 2DGS [17], scale regularization effectively fixes the holes in the ShapeNet [57] dataset.
Figure 8: Impact of the initial number of primitives. We qualitatively demonstrate the impact of the initial number of primitives on the granularity of decomposition. A larger number of initial primitives results in a finer-grained decomposition, whereas a smaller number leads to a coarser decomposition.
Figure 9: Applications. Part-Aware Editing (top): After optimization, we can easily edit the scene by moving, adding, removing, or scaling specific parts. 3D Content Generation (bottom): By combining parts of different objects, we can effortlessly create new and meaningful 3D content.
IV-D Ablations
TABLE VI: Effect of the parsimony loss weight ($\lambda_{\text{par}}$) on DTU [55]. We report Chamfer Distance (CD), rendering metrics, and the number of primitives (#P) of block-level reconstruction and point-level reconstruction.
@@ IV-D Ablations
TABLE VII: Effect of the initial number of primitives ($M$) on DTU [55]. We report Chamfer Distance (CD), rendering metrics, and the number of primitives (#P) of block-level reconstruction and point-level reconstruction.
Block-level Point-level
[…]
@@
In this section, we first verify and analyze the critical strategies proposed by our algorithm: overlapping loss, opacity loss, coverage loss, adaptive number of primitives, and parsimony loss. By removing each strategy one at a time, we report the averaged quantitative performance of reconstruction on 15 instances of chair category in Tab. V and provide visual comparisons in Fig. 6.
As shown in Tab. V, VI, and VII, the point-level reconstruction accuracy and rendering quality do not differ significantly. This is because the refinement stage primarily assists in recovering the details of geometry rather than part decomposition. Therefore, in the subsequent analysis, we will focus on the block-level reconstruction results. Removing the overlapping loss improves reconstruction quality and reduces decomposition semantics while retaining more parts. Intuitively, more superquadrics will better fit a shape, but overlaps diminish the meaningfulness of decomposition, as shown in the first column of Fig. 6. Removing opacity loss drastically affects the number of primitives, reducing them to an average of only 3.8. Simultaneously, the CD increases by approximately twofold. This demonstrates the importance of opacity loss in controlling the existence of primitives and ultimately impacting reconstruction accuracy. In the visual example, the reconstructed chair is missing its legs and armrests. Coverage loss significantly involves reconstruction errors as the CD increases by 38%, ensuring the primitives stay in the correct position to fit the target best. As shown in the third column of Fig. 6, there is space in the chair seat that primitives do not cover without coverage loss. The adaptive number of primitives strategy fills in missing parts and improves the reconstruction metric by 36%. Omitting parsimony loss results in the lowest CD errors and best rendering metrics but produces 0.5 times more primitives than the full model, indicating redundancy and over-decomposition.
@@
we also validate the effectiveness of Gaussian scale regularization in Fig. 7. We find that 2DGS [17] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF. The holes occur since large-scale Gaussians can’t support consistent depth rendering from different views. Scale regularization effectively fixes these holes by preventing the formation of large-scale Gaussians.
@@
Lastly, we analyze the impact of two key hyperparameters of our method on decomposition granularity: the parsimony loss weight $\lambda_{\text{par}}$ and the initial number of primitives $M$. As shown in Tab.VI, $\lambda_{\text{par}}$ directly influences the number of primitives found (#P). Stronger regularization results in fewer primitives, while weaker regularization has the opposite effect. However, having too few primitives severely impacts reconstruction accuracy and rendering quality. In Tab.VII, it can be observed that as $M$ increases, both reconstruction and view synthesis performance slightly improve. Simultaneously, $M$ positively correlates with #P. This indicates that by adjusting the values of $M$ and $\lambda_{\text{par}}$, our approach can control the granularity of object or scene decomposition. Fig. 8 visually represents this impact.
IV-E Applications
@@
Fig. 9 illustrates two applications of our method that original 3DGS and NeRF-based approaches do not support. Firstly, after optimization, we obtain the part decomposition, making it easy to edit specific parts of the object or scene, e.g., adding, moving, removing, or scaling sizes. Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content. By replacing parts, we do not need to adjust the position and part connections manually. For man-made objects, the proposed method supports 3D shape generation within the same categories nicely.
V Conclusion

### v2

@@ 4.3.1 DTU and Shapenet Benchmark
Complete model $2.32$ $24.92$ $0.906$ $0.122$ $\underline{0.94}$ $\underline{36.07}$ $0.977$ $0.038$ $6.6$
w/o $L_{\text{over}}$ $\underline{2.11}$ $\underline{25.17}$ $\underline{0.914}$ $\underline{0.118}$ $1.34$ $36.04$ $\underline{0.977}$ $\underline{0.037}$ $7.6$
w/o $L_{\text{opa}}$ $5.72$ $21.12$ $0.853$ $0.174$ $1.06$ $35.60$ $0.975$ $0.040$ $3.8$
@@
w/o $L_{\text{par}}$ 1.91 25.48 0.918 0.115 0.93 36.18 0.977 0.037 $10.1$
Table 5: Ablation studies on the ShapeNet [6]. We report Chamfer Distance, rendering metrics, and the number of primitives (#P).
4.3.2 Real-life Data
To further demonstrate the applicability of our method for learning shape decomposition, we test our model on the real-life images from BlenderMVS dataset [45] and a self-captured dataset. As shown in Fig. 4, our approach can robustly produce both realistic appearances and meaningful 3D decompositions across a variety of data types. More results are provided in the supplementary material.
Figure 7: Impact of Gaussian scale regularization (SR). The degraded mesh produced by 2DGS [16] is effectively improved.
Figure 8: Impact of the initial number of primitives ($M$). Increasing $M$ yields a finer-grained decomposition, while decreasing it produces a coarser decomposition.
Figure 9: Applications. Part-Aware Editing (top): After optimization, we can easily edit the scene by adding, scaling, or moving specific parts. 3D Content Generation (middle): By combining parts of different objects, we can create new 3D content. Physics simulations (bottom): The representation enables us to perform physics-based simulations easily.
4.3.3 Compared to SAM-based Methods
We also conduct comparisons to SAM-based methods for this task, as shown in Fig. 6. Despite impressive advances in 3D segmentation and editing [5, 19, 46, 47] by distilling 2D information, achieving finer-grained semantic disentanglement remains challenging due to indistinct textures and severe cross-view 2D inconsistencies. Lifting 2D segmentation features to 3D segmentation will also lead to feature misalignment, leading to visual artifacts. In contrast, our method segments directly from 3D space, enabling more fine-grained and clearly defined component decomposition.
4.4 Ablations
@@
Table 6: Effect of parsimony weight ($\lambda_{\text{par}}$) and initial primitives count ($M$) on DTU [17]. We report Chamfer distance, rendering metrics, and primitives count (#P). * denotes the default setting.
In this section, we first analyze the design choices by isolating each strategy to assess its impact. We report the averaged quantitative performance on 15 instances of chair category in Tab. 5 and provide visual comparisons in Fig. 5. Removing overlapping loss enhances reconstruction quality but reduces semantic clarity. Omitting opacity loss drastically affects the number of primitives and doubles the CD, highlighting its role in controlling primitive presence and reconstruction accuracy. Coverage loss is crucial in reconstruction accuracy, ensuring primitives align correctly with the target. The adaptive primitive strategy fills in missing parts and improves the reconstruction metric. Without parsimony loss, CD and rendering metrics are optimal but result in redundancy and over-decomposition.
@@
We also validate the effectiveness of Gaussian scale regularization in Fig. 7. 2DGS [16] tends to use large-scale Gaussians in texture-less areas and create holes in the constructed mesh by TSDF. Holes occur since large-scale Gaussians can’t support consistent depth rendering from different views. Scale regularization effectively addresses this issue by suppressing large-scale Gaussians.
@@
Lastly, in Tab. 6, we analyze the influence of two key hyperparameters on decomposition granularity: the parsimony loss weight $\lambda_{\text{par}}$ and the initial primitive count $M$. Stronger parsimony regularization reduces the number of primitives, while weaker regularization increases it. As $M$ rises, both reconstruction and view synthesis performance slightly improve. This demonstrates that adjusting $M$ and $\lambda_{\text{par}}$ enables our method to effectively control the granularity of object or scene decomposition. Fig. 8 visually illustrates this impact.
4.5 Applications
@@
Fig. 9 illustrates three applications of our method that original 3DGS and NeRF-based approaches do not support. Firstly, after optimization, we obtain the part decomposition, facilitating easy editing of specific object or scene components, e.g., adding, moving, removing, or scaling. Secondly, by combining parts of different objects, our method enables the creation of new high-quality 3D content. Additionally, our method enables easy manipulation of objects within a scene, supporting physics-based simulations.
5 Conclusion